# Iterative Inversion of Ill-Conditioned MIMO Channels Using Symmetric Rank-$1$ Regularization

### v3

@@ V-C5 Experiment 5
We first discuss the performance when the channel is stationary. Fig. 10 shows the average SER with the increase of iteration number when the normalized transmission power is $(\mathcal{E})/(\sigma_{v}^{2})=62.5$ dB for i.i.d. Rayleigh channel and $(\mathcal{E})/(\sigma_{v}^{2})=67.5$ dB for i.i.d. Rician channel, respectively. It is shown that in either case, the e-PIA and PIA converge to RZF performance when $i=36$ and $i=41$, respectively. Among the baselines, the GS converges with the least iteration number ($i=44$) in Rayleigh channel, while the Schulz method alone converges with least iteration number ($i=58$) in Rician channel. This demonstrates the advantage of both the PIA and e-PIA of reducing the convergence iteration number, and coincides with our discussion in Experiment 3.
@@
Notably, GS preconditioning requires more iteration number to converge in Rician channel, although it showed advantage in average condition number in Fig. 6. This is because when the channel is extremely ill-conditioned, GS will has similar behavior as the SSOR (i.e., making the condition number worse). This leads to burst errors when the channel is extremely ill-conditioned, and makes the SER worse.
Fig. 11 shows the case of $M=N=512$, where the SNR is $(\mathcal{E})/(\sigma_{v}^{2})=65$ dB and $(\mathcal{E})/(\sigma_{v}^{2})=70$ dB for the i.i.d. Rayleigh channel and i.i.d. Rician channel, respectively. It is shown that the PIA and e-PIA converges to RZF performance when the iteration number is $i=43$ and $i=46$, respectively. In comparison, the smallest iteration number is $i=48$ of GS preconditioning in Rayleigh channel and $i=65$ of the Schulz method alone in Rician channel. In the case of i.i.d. Rician, the e-PIA significantly reduces the iteration number for $35\%$. This shows that the PIA and e-PIA are able to preserve its advantage as the MIMO size increases.
Figure 10: Average SER of the PIA/e-PIA approaches as the iteration number increases compared to exact RZF performance in i.i.d. Rayleigh channel and i.i.d. Rician channel when $M=N=128$.
Figure 11: Average SER of the PIA/e-PIA approaches as the iteration number increases compared to exact RZF performance in i.i.d. Rayleigh channel and i.i.d. Rician channel when $M=N=512$.
When the channel is spatially non-stationary, the PIA/e-PIA are still able to reduce the iteration number, as shown in Fig. 12. The SNR for the LoS-dominated ELAA is $(\mathcal{E})/(\sigma_{v}^{2})=42.5$ dB and for the mixed LoS/NLoS is $(\mathcal{E})/(\sigma_{v}^{2})=70$ dB. In LoS-dominated ELAA, using the PIA is already close to the performance bound of SR-$1$R, as shown in Fig 8. Hence, we only use the PIA and it converges to RZF performance when $i=19$. In comparison, the GS preconditioning achieves RZF performance when $i=23$. In mixed LoS/non-LoS ELAA, the PIA and e-PIA converges to RZF performance when $i=51$ and $i=48$, respectively. In comparison, the Schulz method alone achieves RZF performance when $i=56$.
Among the linear system inverse approaches, the BFGS algorithm suffers from significant performance degradation. In contrast, the MRIDA always achieves the lowest SER when the iteration number is small, e.g.., $i\leq 10$. This confirms the advantage of the MRIDA in massive-MIMO systems, where the convergence usually occurs within $10$ iterations [9]. This advantage is particularly evident in Fig. 12(a), where the MRIDA achieves the lowest SER when $i\leq 15$.
With the increase of iteration number, the SER of the MRIDA is exceeded by iterative matrix inversion solutions (including using the Schulz method alone, the SR-$1$R and the preconditioning baselines). This phenomenon is not due to lack of convergence, but because of the convergence speed. For the MRIDA, the convergence speed can be reformulated based on [9, Theorem 1]:
$\mathbb{E}(\|\mathbf{G}(\mathbf{x}_{i}-\mathbf{x})\|^{2})\leq\psi^{i}\|\mathbf{G}(\mathbf{x}_{0}-\mathbf{x})\|^{2},$ (86)
where $\mathbb{E}(\cdot)$ stands for the expectation, $\mathbf{G}$ for a weighting matrix, $\mathbf{x}_{i}$ for the output of the $i$-th iteration, $\mathbf{x}_{0}$ for the initial state, $\mathbf{x}$ for the target vector, and $\psi\in(0,1)$ for the convergence rate. The speed of convergence is exponential.
For the iterative matrix inversion, reformulating (II-B2) and replace $\mathbf{A}$ with $\mathbf{R}$ yield $\mathbf{I}-\mathbf{R}\mathbf{X}_{i}=(\mathbf{I}-\mathbf{R}\mathbf{X}_{i-1})^{2}$. Hence, the spectrum radius of $\mathbf{I}-\mathbf{R}\mathbf{X}_{i}$ is given by
$\rho(\mathbf{I}-\mathbf{R}\mathbf{X}_{i})=\rho(\mathbf{I}-\omega\mathbf{R}\mathbf{R}^{H})^{2^{i}}.$ (87)
Since $\omega$ is designed to ensure $\rho(\mathbf{I}-\omega\mathbf{R}\mathbf{R}^{H})<1$, we have $\mathbf{X}_{i}$ converging to $\mathbf{R}^{-1}$ in a local quadratic speed [4]. In ill-conditioned channels, more iterations are required for convergence, where the convergence speed advantage of the Schulz method becomes particularly beneficial.
Figure 12: Average SER of the PIA/e-PIA approaches as the iteration number increases compared to exact RZF performance in LoS-dominated ELAA channel ($M=64$, $N=16$) and mixed LoS/non-LoS ELAA channel ($M=N=128$).
@@
Table I shows the calculated latency. The first thing to be noticed is that the Schulz method is faster than the GS preconditioning in general. This is because of the long computation latency of the GS preconditioning. When $N=128$ or $N=512$, the PIA achieves $2.5\%\sim 28\%$ reduction of latency compared to using the Schulz method alone, while the e-PIA extends this reduction to $5.5\%\sim 32\%$. This demonstrates that the PIA/e-PIA can transform its advantage of fast convergence into shorter computation latency.
@@
V-C6 Experiment 6
While the SR-$1$R has demonstrated great performance improvement, its performance is limited by $\lambda_{1}$ and $\lambda_{N-2}$. It would be interesting to know the potential to go beyond this limit. As requested by the reviewer, here we present an example using a rank-$K$ ($K>1$) matrix to regularize $\mathbf{A}$. In this case, the regularized matrix in (1) becomes:
$\mathbf{R}=\mathbf{A}-\mathbf{B}\mathbf{\Xi}\mathbf{B}^{H},$ (88)
where $\mathbf{B}\in\mathbb{C}^{N\times K}$, and $\mathbf{\Xi}\in\mathbb{C}^{K\times K}$ is a diagonal matrix. In this case, $\mathbf{A}^{-1}$ can be recovered from $\mathbf{R}^{-1}$ with the Woodbury formula [27].
As the theoretical foundation of rank-$K$ regularization is not ready, it is unclear how to determine $\mathbf{B}$ and $\mathbf{\Xi}$ to minimize $\kappa(\mathbf{R})$. Simply repeating the SR-$1$R would not work, as $\theta_{N-1}$ will keep decreasing. After $K$ repetitions, $|\theta_{N-1}|$ will become extremely large and increase the condition number. Therefore, in our example, we specifically look into three representative cases to reveal the potential of the rank-$K$ design.
  • Approach 1: The first approach is to determine $\mathbf{B}$ and $\mathbf{\Xi}$ based on the EVD. Recall (19) where $\mathbf{A}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{H}$. We move the first $K_{1}$ and last $K_{2}$ eigenvalues ($K_{1}+K_{2}=K$) to the average of the rest eigenvalues. In this case, $\mathbf{B}$ and $\mathbf{\Xi}$ are given by
    $\mathbf{B}$ $=[\mathbf{u}_{0},\cdots,\mathbf{u}_{K_{1}-1},\mathbf{u}_{N-K_{2}},\cdots\mathbf{u}_{N-1}],$ (89)
    $\mathbf{\Xi}$ $=\overline{\lambda}_{K_{1},K_{2}}\mathbf{I}-\mathbf{\Lambda}_{K},$ (90)
    where $\mathbf{u}_{n}$ stands for the $n^{th}$ column of $\mathbf{U}$, $\mathbf{\Lambda}_{K}$ for the diagonal matrix formed by the first $K_{1}$ and last $K_{2}$ eigenvalues, and $\overline{\lambda}_{K_{1},K_{2}}$ for the average of other eigenvalues:
    $\overline{\lambda}_{K_{1},K_{2}}=(\lambda_{K_{1}}+\cdots+\lambda_{N-K_{2}-1})/(N-K).$ (91)
  • Approach 2: It is impractical to expect perfect EVD knowledge. Therefore, we add small Gaussian noise to $\mathbf{B}$ and $\mathbf{\Xi}$ to investigate the condition number when the knowledge of $\mathbf{B}$ and $\mathbf{\Xi}$ is not perfect.
  • Performance Bound: The rank-$K$ regularization has been discussed in the literature [47], and we can conclude the following inequalities:
    $\lambda_{n+K}\leq$ $\theta_{n}\leq\lambda_{n},_{~{}n=0,\cdots,N-K-1},$ (92)
    $\lambda_{N-1}-\mathrm{Tr}(\mathbf{\Xi})\leq$ $\theta_{n}\leq\lambda_{n},_{~{}n=N-K,\cdots,N-1}.$ (93)
    Similar to the discussions in Section III-A, (92) and (93) yield a lower bound of the condition number:
    $\kappa(\mathbf{R})\geq\lambda_{K}/\lambda_{N-K-1}.$ (94)
Figure 13: CDF of condition number for the rank-$K$ regularization in i.i.d. Rayleigh channel when $M=N=512$.
In simulation, we consider $\mathbf{H}$ conforming to i.i.d. Rayleigh fading with $M=N=512$. We consider $\mathbf{A}=\mathbf{H}^{H}\mathbf{H}$. For Approach 2, The Gaussian noise added to $\mathbf{B}$ and $\mathbf{\Xi}$ is i.i.d. with a variance of $0.015$. We consider two cases: $K=16$ and $K=8$ where $K_{1}=K_{2}=(K)/(2)$.
Fig. 13 shows the CDF of the condition number after rank-$K$ regularization. As baselines, the green and black lines stand for the SR-$1$R and the original $\mathbf{A}$, respectively. It is shown that Approach 1 reduces the condition number for around $12$ dB when $K=8$. When $K$ increases to $16$, the condition number is reduced for around another $10$ dB. Moreover, surprisingly, Approach 2 outperforms Approach 1 for around $3$ dB both when $K=8$ and $K=16$. This is indeed encouraging and reveals a more optimal design of $\mathbf{B}$ and $\mathbf{\Xi}$ remains to be investigated in the future.
@@ VI Conclusion
In this paper, we have proposed to regularize the condition number of the channel’s Gram matrix with a symmetric rank-$1$ matrix in MIMO systems, termed the SR-$1$R. The SR-$1$R can align the largest/smallest eigenvalues to their closest neighbor when the rank-$1$ matrix is formed by their corresponding eigenvectors. To enable the SR-$1$R in real-time signal processing, we have proposed the PIA approach, where the power iteration is employed to estimate these eigenvectors. Moreover, the e-PIA approach has been proposed to enhance the estimation of the eigenvector for the smallest eigenvalue. Both the PIA/e-PIA approaches have good compatibility to parallel computing with logarithmic algorithm-depth.

### v4

@@ V-C4 Experiment 4
We first discuss the performance when the channel is stationary. Fig. 5(a) and Fig. 5(b) shows the case of $M=N=512$, where the SNR is $(\mathcal{E})/(\sigma_{v}^{2})=65$ dB for the i.i.d. Rayleigh channel and $(\mathcal{E})/(\sigma_{v}^{2})=70$ dB for the i.i.d. Rician channel. It is shown that the PIA and e-PIA converges to RZF performance when the iteration number is $i=43$ and $i=46$, respectively. In comparison, the smallest iteration number is $i=48$ of GS preconditioning in Rayleigh channel and $i=65$ of the Schulz method alone in Rician channel. This yields a reduction of iteration number of $4\%\sim 29\%$ for the PIA and $10\%\sim 35\%$ for the e-PIA, and demonstrates the advantage of both the PIA and e-PIA of reducing the convergence iteration number, coinciding with our discussion in Experiment 1.
@@
Notably, GS preconditioning requires more iteration number to converge in Rician channel, although it showed advantage in average condition number in Fig. 3(c). This is because when the channel is extremely ill-conditioned, GS will has similar behavior as the SSOR (i.e., making the condition number worse). This leads to burst errors when the channel is extremely ill-conditioned, and makes the SER worse.
@@
With the latency measurement in Fig. 3 and the iteration number in Fig. 5, we are able to calculate the overall computation latency of iterative linear precoding, as shown in Table I. We only present the Schulz method alone and the GS preconditioning in Table I since they need the least number of iterations to converge to RZF performance in Fig. 5. The Jacobi preconditioning usually requires the same iteration as the Schulz method alone, while the SSOR requires more iterations than the GS.
Table I shows the calculated latency. The first thing to be noticed is that the Schulz method is faster than the GS preconditioning in general. This is because of the long computation latency of the GS preconditioning. Compared to the Schulz method, the PIA achieves $9.1\%\sim 28\%$ reduction of latency compared to using the Schulz method alone, while the e-PIA extends this reduction to $13.7\%\sim 32\%$. This demonstrates that the PIA/e-PIA can transform its advantage of fast convergence into shorter computation latency.
It is worth noting that the latency results reported in Table I may appear less favorable compared to direct matrix inversion when evaluated on current hardware platforms. This is primarily because existing parallel computing implementations still incurs a latency above linear scalability for large matrices, as demonstrated by the ‘Schulz’ line in Fig. 3. Nonetheless, this limitation is expected to diminish with the continued advancement of parallel processing technologies [5, 6]. In this work, we focus on the comparative performance of iterative methods.
@@ VI Conclusion
In this paper, we have proposed to regularize the condition number of the channel’s Gram matrix with a symmetric rank-$1$ matrix in MIMO systems, termed the SR-11R. The SR-11R can align the largest/smallest eigenvalues to their closest neighbor when the rank-$1$ matrix is formed by their corresponding eigenvectors. To enable the SR-11R in real-time signal processing, we have proposed the PIA approach, where the power iteration is employed to estimate these eigenvectors. Moreover, the e-PIA approach has been proposed to enhance the estimation of the eigenvector for the smallest eigenvalue. Both the PIA/e-PIA approaches have good compatibility to parallel computing with logarithmic algorithm-depth.